Ted Roberts has been offered the following future ​payments n​ years from today. If his opportunity cost is i​, compounded​ annually, what value would he place on each​ opportunity?  

Future

Value​ ($)

Interest

Rate​ (%)

Years

Present  

Value​ ($)

8,700

6

11

5,800

8

25

Future Value $ Interest Rate % Years Present value
6,300 16 27
2,900 12 19

What is the Present Value?
Answer #1

1)

Present value = Future value / (1 + r)n

Present value = 8,700 / (1 + 0.06)11

Present value = 8,700 / 1.898299

Present value = $4,583.05

2)

Present value = Future value / (1 + r)n

Present value = 5,800 / (1 + 0.08)25

Present value = 5,800 / 6.848475

Present value = $846.90

3)

Present value = Future value / (1 + r)n

Present value = 6,300 / (1 + 0.16)27

Present value = 6,300 / 55.000382

Present value = $114.54

4)

Present value = Future value / (1 + r)n

Present value = 2,900 / (1 + 0.12)19

Present value = 2,900 / 8.612762

Present value = $336.71